# No “Zero-Shot” Without Exponential Data: Pretraining Concept Frequency Determines Multimodal Model Performance

### v2

@@ Abstract
Web-crawled pretraining datasets underlie the impressive “zero-shot" evaluation performance of multimodal models, such as CLIP for classification/retrieval and Stable-Diffusion for image generation. However, it is unclear how meaningful the notion of “zero-shot" generalization is for such multimodal models, as it is not known to what extent their pretraining datasets encompass the downstream concepts targeted for during “zero-shot" evaluation. In this work, we ask: How is the performance of multimodal models on downstream concepts influenced by the frequency of these concepts in their pretraining datasets?
We comprehensively investigate this question across 34 models and five standard pretraining datasets (CC-3M, CC-12M, YFCC-15M, LAION-400M, LAION-Aesthetics), generating over 300GB of data artifacts. We consistently find that, far from exhibiting “zero-shot” generalization, multimodal models require exponentially more data to achieve linear improvements in downstream “zero-shot" performance, following a sample inefficient log-linear scaling trend. This trend persists even when controlling for sample-level similarity between pretraining and downstream datasets [79], and testing on purely synthetic data distributions [51]. Furthermore, upon benchmarking models on long-tailed data sampled based on our analysis, we demonstrate that multimodal models across the board perform poorly. We contribute this long-tail test set as the Let it Wag! benchmark to further research in this direction. Taken together, our study reveals an exponential need for training data which implies that the key to “zero-shot” generalization capabilities under large-scale training paradigms remains to be found.
[…]
@@ 1 Introduction
Multimodal models like CLIP [91] and Stable Diffusion [96] have revolutionized performance on downstream tasks—CLIP is now the de-facto standard for “zero-shot" image recognition [133, 72, 126, 48, 132] and image-text retrieval [46, 64, 24, 117, 129], while Stable Diffusion is now the de-facto standard for “zero-shot" text-to-image (T2I) generation [93, 17, 96, 41]. In this work, we investigate this empirical success through the lens of zero-shot generalization [69], which refers to the ability of the model to apply its learned knowledge to new unseen concepts. Accordingly, we ask: Are current multimodal models truly capable of “zero-shot” generalization?
@@
To address this, we conducted a comparative analysis involving two main factors: (1) the performance of models across various downstream tasks and (2) the frequency of test concepts within their pretraining datasets. We compiled a comprehensive list of $4,029$ concepts111class categories for classification tasks, objects in the text captions for retrieval tasks, and objects in the text prompts for generation tasks, see Sec. 2 for more details on how we define concepts. from 27 downstream tasks spanning classification, retrieval, and image generation, assessing the performance against these concepts. Our analysis spanned five large-scale pretraining datasets with different scales, data curation methods and sources (CC-3M [107], CC-12M [27], YFCC-15M [113], LAION-Aesthetics [103], LAION-400M [102]), and evaluated the performance of 10 CLIP models and 24 T2I models, spanning different architectures and parameter scales. We consistently find across all our experiments that, across concepts, the frequency of a concept in the pretraining dataset is a strong predictor of the model’s performance on test examples containing that concept. Notably, model performance scales linearly as the concept frequency in pretraining data grows exponentially i.e., we observe a consistent log-linear scaling trend. We find that this log-linear trend is robust to controlling for correlated factors (similar samples in pretraining and test data [79]) and testing across different concept distributions along with samples generated entirely synthetically [51].
@@
Our findings indicate that the impressive empirical performance of multimodal models like CLIP and Stable Diffusion can be largely attributed to the presence of test concepts within their vast pretraining datasets, thus their reported empirical performance does not constitute “zero-shot” generalization. Quite the contrary, these models require exponentially more data on a concept to linearly improve their performance on tasks pertaining to that concept, highlighting extreme sample inefficiency.
@@
In our analysis, we additionally document the distribution of concepts encountered in pretraining data and find that:
@@
  • Concept Distribution: Across all pretraining datasets, the distribution of concepts is long-tailed (see Fig. 5 in Sec. 5), which indicates that a large fraction of concepts are rare. However, given the extreme sample inefficiency observed, what is rare is not properly learned during multimodal pretraining.
@@
  • Concept Correlation across Pretraining Datasets: The distribution of concepts across different pretraining datasets are strongly correlated (see Tab. 4 in  Sec. 5), which suggests web crawls yield surprisingly similar concept distributions across different pretraining data curation strategies, necessitating explicit rebalancing efforts [11, 125].
@@
  • Image-Text Misalignment between Concepts in Pretraining Data: Concepts often appear in one modality but not the other, which implies significant misalignment (see Tab. 4 in Sec. 5). Our released data artifacts can help image-text alignment efforts at scale by precisely indicating the examples in which modalities misalign. Note that the log-linear trend across both modalities is robust to this misalignment.
@@
To provide a simple benchmark for generalization performance for multimodal models, which controls for the concept frequency in the training set, we introduce a new long-tailed test dataset called “Let It Wag!”. Current models trained on both openly available datasets (e.g., LAION-2B [103], DataComp-1B [46]) and closed-source datasets (e.g., OpenAI-WIT [91], WebLI [29]) have significant drops in performance, providing evidence that our observations may also transfer to closed-source datasets. We publicly release all our data artifacts (over 300GB), amortising the cost of analyzing the pretraining datasets of multimodal foundation models for a more data-centric understanding of the properties of multimodal models in the future.
@@
Several prior works [91, 46, 82, 42, 83, 74] have investigated the role of pretraining data in affecting performance. Mayilvahanan et al. [79] showed that CLIP’s performance is correlated with the similarity between training and test datasets. In other studies on specific areas like question-answering [62] and numerical reasoning [94] in large language models, high train-test set similarity did not fully account for observed performance levels [127]. Our comprehensive analysis of several pretraining image-text datasets significantly adds to this line of work, by (1) showing that concept frequency determines zero-shot performance and (2) pinpointing the exponential need for training data as a fundamental issue for current large-scale multimodal models. We conclude that the key to “zero-shot" generalization capabilities under large-scale training paradigms remains to be found.
@@ 2 Concepts in Pretraining Data and Quantifying Frequency
Figure 1: Concept Extraction and Frequency Estimation Pipeline. (left) We compile $4,029$ concepts from $17$ classification, $2$ retrieval, and $8$ image generation prompt datasets. (right) We construct efficient indices for both text-search (using standard unigram indexing (1)) and image-search (using RAM++ [59] (2)); intersecting hits from both gives us (3) the image-text matched frequencies per concept.
@@
In this section, we outline our methodology for obtaining concept frequencies within pretraining datasets. We first define our concepts of interest, then describe algorithms for extracting their frequencies from images and text captions of pretraining datasets. Finally, we discuss how to aggregate them to calculate matched image-text concept frequencies. For a schematic overview of our methods, see Fig. 1.
@@
Defining Concepts. We define “concepts” as the specific objects or class categories we seek to analyze in the pretraining datasets. For zero-shot classification tasks, these concepts are the class names, such as the $1,000$ classes in ImageNet [35] (e.g., “tench”, “goldfish”, “stingray”). For image-text retrieval and image generation tasks, concepts are identified as all nouns present in the test set captions or generation prompts, respectively. For example, in the caption, “A man is wearing a hat”, we extract “man” and “hat” as relevant concepts. We additionally filter out nouns that are present in less than five downstream evaluation samples to remove ambiguous or irrelevant concepts. Across all our experiments, we collate a list of $4,029$ concepts sourced from $17$ classification, $2$ retrieval, and $8$ image generation downstream datasets (see Tab. 1 for details).
@@
Concept Frequency from Text Captions. To enable efficient concept searches, we pre-index all captions from the pretraining datasets, i.e., construct a mapping from concepts to captions. We first use part-of-speech tagging to isolate common and proper nouns and subsequently lemmatize them to standardize word forms [65] with SpaCy [58] . These lemmatized nouns are then cataloged in inverted unigram dictionaries, with each noun being the key and all the indices in the pretraining data samples containing that noun being its values. To determine the frequency of a concept, particularly those composed of multiple words, we examine the concept’s individual unigrams within these dictionaries. For multi-word expressions, by intersecting the lists of sample indices corresponding to each unigram, we identify the samples that contain all parts of the concept. The frequency of the concept in the text captions is the count of these intersecting sample indices. Our frequency estimation algorithm hence allows scalable $\mathcal{O}(1)$ search with respect to the number of captions for any given concept in the pretraining dataset captions.
@@
Concept Frequency from Images. Unlike text captions, we do not have a finite vocabulary for pre-indexing pretraining images, and thus cannot perform $\mathcal{O}(1)$ concept lookup. Instead, we collect all the $4,029$ downstream concepts and verify their presence in images using a pretrained image tagging model. We tested various open-vocabulary object detectors, image-text matching models and multi-tagging models. We found that RAM++ [59]—an open-set tagging model that tags images based on a predefined list of concepts in a multi-label manner—performs the best. This approach generates a list of pretraining images, each tagged with whether the downstream concepts are present or not, from which we can compute concept frequencies. We provide qualitative examples along with design choice ablations in Appx. F.
@@
Image-Text Matched Concept Frequencies. Finally, we combine the frequencies obtained from both text and image searches to calculate matched image-text frequencies. This involves identifying pretraining samples where both the image and its associated caption correspond to the concept. By intersecting the lists from our image and text searches, we determine the count of samples that align in both modalities, offering a comprehensive view of concept representation across the dataset. We note that this step is necessary as we observed significant image-text misalignment between concepts in the pretraining datasets (see Tab. 4), hence captions may not reflect what is present in the image and vice-versa. This behaviour has also been alluded to in prior work investigating pretraining data curation strategies [76, 75, 124, 83]. We provide more detailed analysis on image-text misalignment in Sec. 5.
3 Comparing Pretraining Frequency & “Zero-Shot" Performance
@@
Having obtained frequency estimates for our downstream concepts, we now establish the relationship between image-text matched pretraining concept frequencies and zero-shot performance across classification, retrieval, and generation tasks. We first detail our experimental approach and then discuss key results.
@@ 3.1 Experimental Setup
3.1.1 Image-Text (CLIP) Models
Datasets. Our evaluation consists of 4 pretraining datasets, 2 downstream retrieval datasets, and 17 downstream classification datasets, presented in Tab. 1, covering a broad spectrum of objects, scenes, and fine-grained distinctions.
Models. We test CLIP [91] models with both ResNet [53] and Vision Transformer [36] architecture, with ViT-B-16 [81] and RN50  [48, 82] trained on CC-3M and CC-12M, ViT-B-16, RN50, and RN101 [61] trained on YFCC-15M, and ViT-B-16, ViT-B-32, and ViT-L-14 trained on LAION400M [102]. We follow open_clip [61], slip [81] and cyclip [48] for all implementation details.
Prompting. For zero-shot classification, we experiment with three prompting strategies: {classname} only, “A photo of a {classname}" and prompt-ensembles [91], which averages over $80$ different prompt variations of {classname}. For retrieval, we use the image or the caption as input corresponding to I2T (image-to-text) or T2I (text-to-image) retrieval respectively.
Metrics. We compute mean zero-shot classification accuracy for classification tasks [91]. For retrieval, we assess performance using traditional metrics for both text-to-image and image-to-text retrieval tasks [91] (Recall@1, Recall@5, Recall@10).
3.1.2 Text-to-Image Models
@@
Datasets. Our pretraining dataset is LAION-Aesthetics [103], with downstream evaluations done on subsampled versions of eight datasets as released by HEIM [71]: CUB200 [121], Daily-DALLE [33], Detection [30], Parti-Prompts [130], DrawBench [98], COCO-Base [73], Relational Understanding [32] and Winoground [114]. Please refer to HEIM [71] for more details on the evaluation datasets used.
@@
Metrics. Evaluation consists of image-text alignment and aesthetic scores. For automated metrics [71], we use expected and max CLIP-score [57] to measure image-text alignment along with expected and max aesthetics-score [102] to measure aesthetics. To verify reliability of the automated metrics, we compare them with human-rated scores (measured on a 5-point grading scale) for both image-text alignment and aesthetics [71]. To supplement the human-rated scores provided by HEIM [71], we confirm our findings by performing a small-scale human evaluation as well (see Appx. C).
@@
Figure 2: Log-linear relationships between concept frequency and CLIP zero-shot performance. Across all tested architectures (RN50, RN101, ViT-B-32, ViT-B-16, ViT-L-14) and pretraining datasets (CC-3M, CC-12M, YFCC-15M, LAION-400M), we observe a consistent linear relationship between CLIP’s zero-shot performance on a concept and the log-scaled concept pretraining frequency. This trend holds for both zero-shot classification (results averaged across 17 datasets) and image-text retrieval (results averaged across 2 datasets). ** indicates that the result is significant ($p<0.05$ with a two-tailed t-test.), and thus we show pearson correlation ($\rho$) as well.
@@
Figure 3: Log-linear relationships between concept frequency and T2I aesthetic scores. Across all tested T2I models pretrained on the LAION-Aesthetics dataset, we observe a consistent linear relationship between zero-shot performance on a concept and the log-scaled concept pretraining frequency.
@@
3.2 Result: Pretraining Frequency is Predictive of “Zero-Shot" Performance

### v3

@@ Abstract
Web-crawled datasets underlie the impressive “zero-shot” performance of multimodal models, such as CLIP for classification and Stable-Diffusion for image generation. However, it is unclear how meaningful the notion of “zero-shot” generalization is for such models because the extent to which their pretraining datasets encompass downstream concepts used in “zero-shot” evaluation is unknown. In this work, we ask: How is the performance of multimodal models on downstream concepts influenced by the frequency of these concepts in their pretraining datasets?
We comprehensively investigate this question across 34 models and 5 standard pretraining datasets, generating over 300GB of data artifacts. We consistently find that, far from exhibiting “zero-shot” generalization, multimodal models require exponentially more data to achieve linear improvements in downstream “zero-shot” performance, following a sample inefficient log-linear scaling trend. This trend persists even when controlling for sample-level similarity between pretraining and evaluation datasets [81], and testing on purely synthetic data distributions [52]. Furthermore, upon benchmarking models on long-tailed data sampled based on our analysis, we demonstrate that multimodal models across the board perform poorly. We contribute this long-tail test dataset as the Let it Wag! benchmark to further research in this direction. Taken together, our study reveals an exponential need for training data which implies that the key to “zero-shot” generalization capabilities under large-scale training data and compute paradigms remains to be found.
@@ 1 Introduction
Multimodal models like CLIP [98] and Stable Diffusion [104] have revolutionized performance on downstream tasks. CLIP is now the de facto standard for “zero-shot” image recognition [143, 74, 136, 49, 142] and image-text retrieval [47, 64, 25, 127, 139], while Stable Diffusion is now the de facto standard for “zero-shot” text-to-image (T2I) generation [100, 18, 104, 42]. In this work, we investigate this empirical success through the lens of zero-shot generalization [70], which refers to the ability of models to apply their learned knowledge to new unseen concepts (not seen during training). Accordingly, we ask: Are current multimodal models truly capable of “zero-shot” generalization?
@@
To tackle this question, we conduct a comparative analysis involving two main factors: (1) the performance of models across various downstream tasks, and (2) the frequency of test concepts within their pretraining datasets. We compile a comprehensive list of $4,029$ concepts111class categories for classification tasks, objects in the text captions for retrieval tasks, and objects in the text prompts for generation tasks, see Sec. 2 for more details on how we define concepts. from 27 downstream tasks spanning classification, retrieval, and image generation, assessing model performance against these concepts. Our analysis spanned five large-scale image-text pretraining datasets with different scales, data curation methods and sources (CC-3M [115], CC-12M [28], YFCC-15M [123], LAION-Aesthetics [111], LAION-400M [110]), and evaluated the performance of 10 CLIP models and 24 T2I models, spanning different architectures and parameter scales. We consistently find across our experiments that, across concepts, the frequency of a concept in the pretraining dataset is a strong predictor of the model’s performance on test examples containing that concept (see Fig. 2). Notably, model performance scales linearly as the concept frequency in pretraining data grows exponentially i.e., we observe a consistent log-linear scaling trend. We find that this log-linear trend is robust to controlling for correlated factors (similar samples in pretraining and test data [81]) and testing across different concept distributions along with samples generated entirely synthetically [52].
@@
Our findings indicate that the impressive empirical performance of multimodal models like CLIP and Stable Diffusion can be largely attributed to the presence of test concepts within their vast pretraining datasets, thus their reported empirical performance does not constitute “zero-shot” generalization. Quite the contrary, these models require exponentially more data on a concept to linearly improve their performance on tasks pertaining to that concept, suggesting significant sample inefficiency.
@@
We additionally document the distribution of concepts encountered in pretraining data and find that:
@@
  • Concept Distribution: Across all pretraining datasets, the distribution of concepts is long-tailed (see Fig. 5), indicating that a large fraction of concepts are rare. Given the extreme sample inefficiency observed, these rare concepts are not properly learned during pretraining.
@@
  • Concept Correlation across Pretraining Datasets: The distributions of concepts across different pretraining datasets are strongly correlated (see Tab. 4), suggesting that web crawls yield surprisingly similar concept distributions across very diverse data curation strategies. This necessitates explicit concept rebalancing efforts explored in prior work [11, 135].
@@
  • Image-Text Misalignment in Pretraining Data: Concepts often appear in one modality but not the other, implying significant misalignment (see Tab. 4). Our released data artifacts can help image-text alignment efforts at scale by precisely indicating examples where modalities misalign. Note that the log-linear trend across both modalities is robust to this misalignment.
@@
To provide a simple benchmark for multimodal generalization that controls for concept frequency in the pretraining set, we introduce a new long-tailed test set, “Let It Wag!”. Current models trained on both openly available datasets (e.g., LAION-2B [111], DataComp-1B [47]) and closed-source datasets (e.g., OpenAI-WIT [98], WebLI [30]) have significant drops in performance (see Fig. 6), suggesting that our findings may also transfer to closed-source datasets. We publicly release all data artifacts, amortising the cost of analyzing image-text pretraining datasets for future efforts focused on a more data-centric understanding of the properties of multimodal models.
@@
Situating our Contributions in Broader Literature. Our comprehensive analysis of several image-text datasets significantly adds to prior investigations on the role of pretraining data in affecting performance for both CLIP [92, 81, 43] and language models [62, 102, 82], by (1) showing that concept frequency determines zero-shot performance, and (2) pinpointing the exponential need for training data as a fundamental issue for current multimodal foundation models. We conclude that the key to “zero-shot” generalization under large-scale training paradigms remains to be found.
@@ 2 Concepts in Pretraining Data and Quantifying Frequency
Figure 1: Concept Extraction and Frequency Estimation. (left) We compile $4,029$ concepts from $27$ evaluation datasets. (right) We construct efficient indices for text-search (unigram indexing (1)) and image-search (RAM++ (2)); intersecting hits from both gives (3) image-text matched frequencies.
@@
In this section, we discuss how to estimate concept frequencies within pretraining datasets. We first define our concepts of interest, then describe algorithms for extracting their individual frequencies from images and text captions of pretraining datasets independently, and describe how we aggregate them to compute matched image-text concept frequencies. For a schematic overview, see Fig. 1.
@@
Defining Concepts. We define “concepts” as the specific objects/relations we seek to analyze in pretraining datasets. Since our goal is to analyze downstream performance of models, we source concepts from 27 target evaluation datasets. For zero-shot classification datasets, extracted concepts are class names, such as the $1,000$ object classes in ImageNet [36] (e.g., tench, goldfish). We also include relational verbs and verb-noun combinations since they are the classes of the UCF101 dataset [116] (e.g., diving, brushing teeth) as well as background nouns from the SUN397 dataset [133] (e.g., abbey, sky). For retrieval and image generation datasets, concepts are all nouns in test set captions or generation prompts. For example, from, “A man is wearing a hat”, we extract “man” and “hat” as concepts. We filter out ambiguous or irrelevant concepts that are present in less than five downstream evaluation samples. In sum, we collate $4,029$ concepts sourced from $17$ classification, $2$ image-text retrieval, and $8$ text-to-image generation downstream datasets (see Tab. 1 and Sec. 3.1 for details).
@@
Concept Frequency from Captions. For efficient concept searches, we pre-index all captions from the pretraining datasets, i.e., construct a mapping from concepts to captions. We first use part-of-speech tagging to isolate common and proper nouns, and lemmatize them with SpaCy [58] (lemmatization helps standardize verbs, enabling the estimation of their frequencies too [65]). These lemmatized terms are then cataloged in inverted unigram dictionaries, mapping each term to all sample indices in the pretraining dataset containing that term. To determine the frequency of a concept, we examine the concept’s unigrams within these dictionaries. For multi-word concepts, we split them into their constituent unigrams, and then independently search for all unigrams before intersecting their hit lists to get a list of matched sample indices. The frequency of the concept in text captions is the count of these intersecting sample indices. This algorithm hence allows scalable $\mathcal{O}(1)$ search with respect to the number of captions for any concept in pretraining dataset captions.
@@
Concept Frequency from Images. Unlike text captions, we do not have a finite vocabulary for pre-indexing pretraining images. Instead, we collect all the $4,029$ downstream concepts and verify their presence in images using a pretrained image tagging model. We tested various open-vocabulary object detectors, multi-tagging models and image-text matching models, for this concept tagging task. We found that RAM++ [59]—an open-set tagging model that tags images based on a predefined list of concept descriptions, in a multi-label manner—performs the best. We automatically consider the relationship between concepts (like synonyms) and concept hierarchies [84], since RAM++ uses descriptions generated by a language model (Sec. I.3) for each concept, to tag each image with certain concepts. This approach generates a list of pretraining images, each tagged with whether the downstream concepts are present or not, from which we can compute concept frequencies (Appx. I).
@@
Image-Text Matched Concept Frequencies. Finally, we combine the frequencies obtained from both text and image searches to compute matched image-text frequencies. This involves identifying pretraining samples where both the image and its associated caption correspond to the concept. By intersecting the lists from our image and text searches, we determine the count of samples that align in both modalities, offering a comprehensive view of concept representation in the pretraining datasets. This step is necessary as we observed significant image-text misalignment between concepts in pretraining datasets (see Tab. 4), hence captions may not reflect what is present in the image and vice-versa. This behaviour has also been alluded to in prior work investigating data curation strategies [78, 77, 134, 89]. We provide a more detailed analysis of image-text misalignment in Sec. 5.
3 Comparing Pretraining Frequency & “Zero-Shot” Performance
@@
Equipped with frequency estimates for downstream concepts, we now establish the relationship between image-text-matched pretraining concept frequencies and zero-shot performance across classification, retrieval, and generation tasks. We first detail our setup and then discuss key results.
@@ 3.1 Experimental Setup
3.1.1 Image-Text (CLIP) Models
Datasets. We use 4 pretraining, 2 downstream retrieval, and 17 downstream classification datasets, covering a broad spectrum of objects, scenes, camera-types, and fine-grained distinctions (see Tab. 1).
Note on Pretraining Dataset Diversity. Each analyzed pretraining dataset significantly differs in data collection, filtering, and cleaning operations. CC-3M [115], originally intended to be used for training image captioning models, explicitly has no real-world entities or proper nouns present, and is cleaned only to have common nouns in its captions. CC-12M [28] and YFCC-15M [123], collected from Flickr, have user-provided metadata. Finally, LAION-400M [110] and LAION-Aesthetics [111] contain raw images downloaded from Common-Crawl [101] with alt-texts as the captions, which can be inherently very noisy as they are uploaded by non-expert humans as a placeholder for images.
Models. We test CLIP [98] models with both ResNet [54] and Vision Transformer [37] architecture, with ViT-B-16 [87] and RN50 [49, 88] trained on CC-3M and CC-12M, ViT-B-16, RN50, and RN101 [61] trained on YFCC-15M, and ViT-B-16, ViT-B-32, and ViT-L-14 trained on LAION400M [110]. We follow open_clip [61], slip [87] and cyclip [49] for our implementation.
Prompting. For zero-shot classification, we experiment with three prompting strategies: {classname} only, “A photo of a {classname}” and prompt-ensembles [98], which averages over $80$ different prompt variations of {classname}. For retrieval, we use the image or the caption as input corresponding to I2T (image-to-text) or T2I (text-to-image) retrieval respectively.
Metrics. We compute mean accuracy for classification tasks [98]. For retrieval, we measure Recall@1, Recall@5, and Recall@10 for both text-to-image and image-to-text retrieval tasks [98].
3.1.2 Text-to-Image Models
Datasets. Our pretraining dataset is LAION-Aesthetics [111], with downstream evaluations done on subsampled versions of eight datasets: CUB200 [131], Daily-DALLE [34], Detection [31], Parti-Prompts [140], DrawBench [106], COCO-Base [75], Relational Understanding [33] and Winoground [124]. Please refer to HEIM [72] for more details on the evaluation datasets.
@@ 3.1.2 Text-to-Image Models
Metrics. Evaluation consists of image-text alignment and aesthetic scores. For automated metrics [72], we use expected and max CLIP-score [57] to measure image-text alignment along with expected and max aesthetics-score [110] to measure aesthetics. To verify reliability of automated metrics, we compare them with human-rated scores (measured on a 5-point grading scale) for both image-text alignment and aesthetics [72]. To supplement the human-rated scores provided by HEIM [72], we confirm our findings by performing our own small-scale human evaluation (Appx. C).
@@
Figure 2: Log-linear relationships between concept frequency and CLIP zero-shot performance. Across all tested architectures (RN50, RN101, ViT-B-32, ViT-B-16, ViT-L-14) and pretraining datasets (CC-3M, CC-12M, YFCC-15M, LAION-400M), we observe a consistent linear relationship between CLIP’s zero-shot performance on a concept and the log-scaled pretraining concept frequency. This trend holds for both zero-shot classification (results averaged across 17 datasets) and image-text retrieval (results averaged across 2 datasets). ** indicates that the result is significant ($p<0.05$ with a two-tailed t-test [118]), and thus we show Pearson correlation ($\rho$[73] as well.
@@
Figure 3: Log-linear relationships between concept frequency and T2I aesthetic scores. Across all tested T2I models pretrained on LAION-Aesthetics, we observe a consistent linear relationship between aesthetic score (averaged across 8 datasets) on a concept and the log-scaled concept frequency.
@@
3.2 Result: Pretraining Concept Frequency is Predictive of “Zero-Shot” Performance